# Addressing Loss of Plasticity and Catas- trophic Forgetting in Continual Learning

### v1

@@ Appendix C Approximated Feature Utility
When we use an origin-passing activation function, $\sigma(0)=0$, we can instead compute the feature utility by setting its input to zero. Thus, the feature utility can be approximated by expanding the loss around the current activation input $a_{l,i}$ as follows:
@@ Appendix F The GetDerivatives and GateDerivatives Functions
Compute the predicted probability vector ${\bm{q}}\leftarrow\bm{\sigma}({\bm{a}}_{L})$ and Compute the loss $\mathcal{L}({\bm{p}},{\bm{q}})$.
@@ Appendix H UPGD on Non-Stationary Toy Regression Problem
In this problem, the task is to calculate the average of two inputs or its negative out of 16 inputs. We introduce non-stationarity using two ways: changing the multiplier $a$ or changing the input-index set $\mathcal{S}$. The learner is required to match the targets by minimizing the online squared error. The learner uses a multi-layer ($300\times 150$) linear network, where the activation used is the identity activation ($\sigma({\bm{x}})={\bm{x}}$). We use linear activations to see if catastrophic forgetting and loss of plasticity may occur even in such simple networks.
@@ Appendix I Experimental Details
In the problem of Label-permuted EMNIST, the labels are permuted every $2500$ sample. For the offline variation, we form a held-out set comprised of $250$ samples of the test set of EMNIST. The performance is measured based on the learner’s prediction on this held-out set after the end of each task in an offline manner. We permute the labels of the held-out set with the same permutation in the task presented to the learner to have consistent evaluation. We observe the same phenomenon indicated by Lesort et al. (2023), in which the performance of SGD decreases on the first few tasks, matching the vast majority of works on catastrophic forgetting. However, when scaling the number of tasks, SGD recovers and maintains its performance. The online variation uses the same experimental details in Section 4.4. The hyper-parameters used are given in Table 3.
@@ I.6 Policy Collapse Experiment
$\beta_{w}$$\{0.9,0.99,0.999\}$ S-EWC, S-SI, S-MAS, S-RWalk
@@
UPGD-W $\alpha=0.01,\sigma=0.001,\beta_{u}=0.999,\lambda=0.0$
N-UPGD-W $\alpha=0.01,\sigma=0.01,\beta_{u}=0.99,\lambda=0.001$
@@
Label-permuted S-EWC $\alpha=0.01,\beta_{I}=0.999,\beta_{w}=0.999,\kappa=1.0$
EMNIST S-MAS $\alpha=0.01,\beta_{I}=0.999,\beta_{w}=0.999,\kappa=10.0$
@@
S-RWalk $\alpha=0.01,\beta_{I}=0.9,\beta_{w}=0.999,\kappa=0.1$
@@
N-UPGD-W $\alpha=0.01,\sigma=0.01,\beta_{u}=0.999,\lambda=0.001$
@@
Label-permuted S-EWC $\alpha=0.01,\beta_{I}=0.999,\beta_{w}=0.999,\kappa=1.0$
mini-ImageNet S-MAS $\alpha=0.01,\beta_{I}=0.999,\beta_{w}=0.999,\kappa=10.0$
@@
S-RWalk $\alpha=0.01,\beta_{I}=0.99,\beta_{w}=0.999,\kappa=0.1$
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
The generate-and-test method (Mahmood & Sutton 2013) is a method that finds better features using search, which, when combined with gradient descent (see Dohare et al. 2023a), is similar to a feature-wise variation of our method. However, this method only works with networks with single-hidden layers in single-output regression problems. It uses the weight magnitude to determine feature utilities; however, it has been shown that weight magnitude is not suitable for other problems, such as classification (Elsayed 2022). On the contrary, our variation uses a better notion of utility that enables better search in the feature space and works with arbitrary network structures or objective functions so that it can be seen as a generalization of the generate-and-test method.

### v2

@@ Appendix C Approximated Feature Utility
When we use an origin-passing activation function, $\sigma_{a}(0)=0$, we can instead compute the feature utility by setting its input to zero. Thus, the feature utility can be approximated by expanding the loss around the current activation input $a_{l,i}$ as follows:
@@ Appendix F The GetDerivatives and GateDerivatives Functions
Compute the predicted probability vector ${\bm{q}}\leftarrow\bm{\sigma}_{a}({\bm{a}}_{L})$ and Compute the loss $\mathcal{L}({\bm{p}},{\bm{q}})$.
@@ Appendix H UPGD on Non-Stationary Toy Regression Problem
In this problem, the task is to calculate the average of two inputs or its negative out of 16 inputs. We introduce non-stationarity using two ways: changing the multiplier $a$ or changing the input-index set $\mathcal{S}$. The learner is required to match the targets by minimizing the online squared error. The learner uses a multi-layer ($300\times 150$) linear network, where the activation used is the identity activation ($\bm{\sigma}_{a}({\bm{x}})={\bm{x}}$). We use linear activations to see if catastrophic forgetting and loss of plasticity may occur even in such simple networks.
@@ Appendix I Experimental Details
In the problem of Label-permuted EMNIST, the labels are permuted every $2500$ sample. For the offline variation, we form a held-out set comprised of $250$ samples of the test set sampled after the end of each task. The performance is measured based on the learner’s prediction on this held-out set after the end of each task in an offline manner. We permute the labels of the held-out set with the same permutation in the task presented to the learner to have consistent evaluation. We observe the same phenomenon indicated by Lesort et al. (2023), in which the performance of SGD decreases on the first few tasks, matching the vast majority of works on catastrophic forgetting. However, when scaling the number of tasks, SGD recovers and maintains its performance. The online variation uses the same experimental details in Section 4.4. The hyper-parameters used are given in Table 3.
@@ I.6 Policy Collapse Experiment
$\beta_{w}$$\{0.9,0.99,0.999\}$ for Input-permuted MNIST S-EWC, S-SI, S-MAS, S-RWalk
and $\{0.9,0.99,0.999,0.9999\}$ for the rest
@@
UPGD-W $\alpha=0.01,\sigma=0.0,\beta_{u}=0.999,\lambda=0.0001$
N-UPGD-W $\alpha=0.01,\sigma=0.0,\beta_{u}=0.999,\lambda=0.001$
@@
Label-permuted S-EWC $\alpha=0.01,\beta_{I}=0.9999,\beta_{w}=0.9999,\kappa=1.0$
EMNIST S-MAS $\alpha=0.01,\beta_{I}=0.999,\beta_{w}=0.9999,\kappa=1.0$
@@
S-RWalk $\alpha=0.01,\beta_{I}=0.99,\beta_{w}=0.9999,\kappa=0.1$
@@
N-UPGD-W $\alpha=0.01,\sigma=0.01,\beta_{u}=0.9999,\lambda=0.001$
@@
Label-permuted S-EWC $\alpha=0.01,\beta_{I}=0.999,\beta_{w}=0.9999,\kappa=1.0$
mini-ImageNet S-MAS $\alpha=0.01,\beta_{I}=0.999,\beta_{w}=0.9999,\kappa=1.0$
@@
S-RWalk $\alpha=0.01,\beta_{I}=0.9,\beta_{w}=0.9999,\kappa=0.1$
@@
UPGD can be seen as an extension of prior ideas that suggest remedying some of the pathological issues in neural networks caused by gradient descent using a non-gradient-based approach called generate-and-test (e.g., Dohare et al. 2023a, Mahmood 2017, Mahmood & Sutton 2013). These methods use search to alter the features based on their usefulness. The hints of these ideas can be traced back to much earlier works (e.g., Kaebling 1994, Lecun et al. 1989, Sutton 1986, Selfridge 1958). Mahmood and Sutton (2013) devised a generate-and-test method that searches for better features of networks with a single hidden layer. Dohare et al. (2023a) extended the method to multiple layers and combined it with backpropagation. Both methods utilize the weight magnitude to determine feature utilities; however, it has been shown that weight magnitude may not be suitable for other problems, such as classification (Elsayed 2022). Moreover, these prior works did not address catastrophic forgetting. UPGD addresses both issues, utilizes a better notion of utility that enables better search in the feature space, and works with arbitrary network structures or objective functions. Hence, UPGD can be seen as a generalization of the prior works on generate-and-test.